Engineers must consider the breadths of male heads when designing helmets. The company researchers have determined that the population of potential clientele have head breadths that are normally distributed with a mean of 6.7-in and a standard deviation of 1.1-in. Due to financial constraints, the helmets will be designed to fit all men except those with head breadths that are in the smallest 1.4% or largest 1.4%. What is the minimum head breadth that will fit the clientele? min = What is the maximum head breadth that will fit the clientele? min =
Answer #1

Given that,

mean = \mu = 6.7

standard deviation = \sigma = 1.1

(A)Using standard normal table,

P(Z < z) =1.4 %

= P(Z < z) = 0.014  

= P(Z <-2.20 ) = 0.014

z = -2.20

Using z-score formula  

x= z * \sigma + \mu

x= -2.20 *1.1+6.7

x= 4.28

minimum = 4.28

(B)Using standard normal table,

P(Z > z) = 1.4%

= 1 - P(Z < z) = 0. 014

= P(Z < z ) = 1 - 0.014

= P(Z < z ) = 0.986

= P(Z < 2.20) = 0. 986

z = 2.20

Using z-score formula  

x = z \sigma + \mu

x = 2.20 *1.1+6.7

x = 9.12

maximum= 9.12